Suppose that there are 250 identical individual consumers in the market for commodity X, each with a demand function given by dx = 6-Px and 50 identical producers of commodity X, each with a supply function given by sx =5Px. (8 Marks)

a)     Derive the market demand function and the market supply function for commodity X.

b)     Compute the market equilibrium price and equilibrium quantity mathematically.

c)     Tabulate the market demand schedule and the market supply schedule for commodity X.

d)    show whether surplus or shortage occurs at 

i)                   P = 2

ii)                 P = 5


1
Expert's answer
2021-07-08T15:26:38-0400

a.

"Q_d=(6-P_x)250\\=1500-250P_x"


"Q_s=(5P_x)50\\=250P_x"


b.

"Q_d=Q_s\\1500-250P_x=250P_x\\1500=500P_x\\P_x=\frac{1500}{500}\\P_x=3"


"Q_d=1500-250(3)=750\\Q_s=250(3)=750"


equilibrium quantity is 750

equilibrium price is 3


c.



d.

At P=2

There is shortage since the quantity demanded exceeds the quantity supplied.

"QS=500\\

Qd=100\\

deficit=500-1000\\

=-500"


At P=5

There is surplus because the quantity supplied exceeds the quantity demanded

"Qd=250\\Qs=1250\\surplus=1250-250\\=1000"